[figure: 9.7 cm]

IMPACT OF RADIATION DEFECTS AND CHEMICAL SUBSTITUTIONS ON THE FLUX PINNING IN Hg-BASED

SUPERCONDUCTING SINGLE CRYSTALS

A. Wiśniewski^{a}, R. Puźniak^{a}, R. Szymczak^{a}, J. Karpiński^{b} a Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warsaw, POLAND b Laboratorium für Festkörperphysik ETH, CH-8093 Zürich, SWITZERLAND

We have compared and contrasted the vortex state properties of single crystals of (Hg,M)Ba_{2}Ca_{n−1}Cu_{n}O_{2n+2+x} with n = 1 or 3, M = Pb or Re, and different oxygen content. Magnetization measurements were performed for magnetic field perpendicular to CuO_{2} planes. The influence of extended neutron defects for vortex pinning also was determined. We found that flux pinning properties were essentially unchanged by partial replacement of Hg with Re or Pb. On the other hand, introduction of radiation defects changed significantly both irreversible and reversible properties of vortex state. The character of the irreversibility line (IL) changed, giving evidence that the system of ”pancake vortices” became more three-dimensional. The significant reduction in the superconductive mass anisotropy parameter γ was observed after irradiation. It was found that the oxygen content is very relevant for flux pinning properties, affecting not only T_{c} but also anisotropy parameter γ. It was shown that the position of the IL depends very strongly on γ.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : A. Wiśniewski

Address for correspondence : Dr A. Wiśniewski Instytut Fizyki PAN Al. Lotników 32/46 02-668 Warszawa

Email address : wisni@ifpan.edu.pl